Aqueous hydrobromic acid HBr reacts with solid sodium hydroxide NaOH to produce aqueous sodium bromide NaBr and liquid water H2O. What is the theoretical yield of water formed from the reaction of 1.6g of hydrobromic acid and 1.1g of sodium hydroxide? Round your answer to 2 significant figures.
Answer #1

The molecular weight of HBr is 80.91g

The number of moles of HBr = 1.6/80.91

= 0.01971

The molecular weight of NaOH = 40g

The number of moles of NaOH= 1.1/40

= 0.0275

NaOH+ HBr = NaBr + H2O

1 mole of HBr formes 1moles of H2O

0.01971 moles of HBr formes =( 1 mole of water /1moles of HBr )×0.01971 mole of HBr

=0.01971 mole of water

1mol of water = 18 g of water

0.01971 mole of water= 18×0.01971

=0.3547g ( theoretical yield)